Please help me to solve  (d^2 y)/(dx^2 )+dy/dx+y=e^(-x).

Grade:12th pass

1 Answers

Deepak Kumar Shringi
5 years ago
y = e-x
dy/dx = -e-x
f’’(x)= e-x